OpenStudy (anonymous):

How do you solve this Luis and Julia have the same number of coins. Luis has only dimes and Julia has only quarters. If Julia has $1.80 more than Luis does, how many coins does each have?

OpenStudy (anonymous):

Too hard?

OpenStudy (anonymous):

Hello???

OpenStudy (anonymous):

j +1.80 = l .25j = .10l

OpenStudy (anonymous):

sorry my son interrupted me

OpenStudy (anonymous):

j = L -1.80 .25 (L-1.80) = .1L .25L -.45 = .1L .25L -.10L = .45 .15L = .45 L = 3

OpenStudy (anonymous):

Ok I got, thanks

OpenStudy (anonymous):

.10(3) = .25J .30 = .25J 1.2 = J